I didnt get how to write an equationin in this problem; Write an equation of a circle with center at the origin and radiu 6.
Could someone help me with this??

Found 3 solutions by stanbon, Alan3354, scott8148:
Answer by stanbon(48535)

I didnt get how to write an equationin in this problem; Write an equation of a circle with center at the origin and radiu 6.
-----------------------------
If center is at (0,0) and radius is "r", the equation is
x^2 + y^2 = r^2
----------------------
Since r = 6 you get:
x^2 + y^2 = 36
=====================
Cheers,
Stan H.

Answer by Alan3354(21583)

Standard circle eqn is
(x-h)^2 + y-k)^2 = r^2
(h,k) is the center, for this problem they're both 0.
--> x^2 + y^2 = 36

Answer by scott8148(5880)

the general equation for a circle centered at (h,k) with a radius of r is (x-h)^2 + (y-k)^2 = r^2

in this case, h and k are both zero (at the origin)
___ so the equation is x^2 + y^2 = 36